A straight wire of mass 300g and length 25m and current of 35A.It is suspended in midair by a uniform horizontal magnetic field B.What is the magnitude of the magnetic field?
1
Expert's answer
2020-12-17T04:52:05-0500

L = 25 m

m = 300 g = 0.3 kg

I = 35 A

The force acting on the current carrying wire in uniform magnetic field

F = BILsinθ

F= BIL (θ = 90º)

Weight of the wire w = mg

BIL = mg

"B = \frac{mg}{IL} \\

B = \frac{0.3 \times 9.8}{35 \times 25} = 0.00336 = 3.36 \times 10^{-3} \;T"

Answer: "3.36 \times 10^{-3} \;T"